Use the Bohr model to address this question. When a hydrogen atom makss a transitlon from the 5h energy level to the 2, counting the ground level as the first, what is the energy of the emitted photon? Express the answer in electron volts. Number eV What is the wavelength, in nanometers, of the emitted photon? Number nm What is the radius, in nanometers, of the hydrogen atom in its initial (5t) energy level? Number nm
Answer #1

energy in nth state is given by E=-13.6/n2 eV

Energy emitted in transitioning from 5th to 2nd state is = -13.6/52 + 13.6/22 = 2.856 eV

Energy emitted in Joules = 2.856*1.6*10-19 = 4.5696*10-19 Joule

We know, E=hc/\lambda

\lambda=hc/E = (6.626*10-34 * 3*108)/4.5696*10-19 =4.35*10-7 m = 435 nm